OpenStudy (anonymous):

What value is added to both sides of the equation x2 - 8x = -10 in order to solve by completing the square?

OpenStudy (asnaseer):

first - do you know what "completing the square" means?

OpenStudy (anonymous):

noo lol thats why I am so confused

OpenStudy (asnaseer):

ok, what the question is asking you to do is to write the left-hand-side of the expression in such a way that you end up with something like:\[(x-a)^2\]where 'a' is some number

OpenStudy (anonymous):

Ok how do I get there lol like what are the steps

OpenStudy (asnaseer):

ok, so first can you please expand the expression \((x-a)^2\)

OpenStudy (asnaseer):

do you know how to do that?

OpenStudy (anonymous):

Ok so the answers they give you a. -8 b. 16 c. -4 d. 4 When i tried doing it out I got 16 is that correct

OpenStudy (asnaseer):

do you just want the answer or do you want to learn how to do this yourself?

OpenStudy (anonymous):

I looked it up in a book and I understand it know more but not sure if I still did it right

OpenStudy (asnaseer):

so lets go through the steps, first can you expand \((x-a)^2\) ?

OpenStudy (anonymous):

I don't understand what you mean

OpenStudy (asnaseer):

do you know how to expand the expression \((x-a)^2\) ?

OpenStudy (anonymous):

What did was I remembered I had to take -8 divide it by 2 then square the answer you get and add it to both sides

OpenStudy (asnaseer):

thats a very /mechanical/ way of doing it. if you follow through the steps with me then you will end up /understanding/ how to solve problems like this.

OpenStudy (anonymous):

okk

OpenStudy (asnaseer):

so, back to our expression, do you know how to expand \((x-a)^2\) ?. i.e. do you know how to expand this:\[(x-a)^2=(x-a)(x-a)=???\]

OpenStudy (anonymous):

ya

OpenStudy (asnaseer):

what do you get?

OpenStudy (anonymous):

That you have to expand it x-a^2 he (x-a)(x-a)

OpenStudy (asnaseer):

\[(x-a)^2=(x-a)(x-a)=x^2-...\]do you know what the rest of the terms on the right-hand-side should be?

OpenStudy (anonymous):

noo

OpenStudy (asnaseer):

then you will have a lot of trouble understanding how to solve these types of problems.

OpenStudy (asnaseer):

let me show you what to do this time and hopefully you will learn for next time.

OpenStudy (asnaseer):

so, we have:\[(x-a)^2=(x-a)(x-a)=x(x-a)-a(x-a)=x^2-ax-ax+a^2\]\[\qquad=x^2-2ax+a^2\]

OpenStudy (asnaseer):

now, if we compare this to the left-hand-side of the question you were given we see you have:\[x^2-8x=...\]

OpenStudy (asnaseer):

so compare that to:\[x^2-2ax+a^2\]and you will notice that 2a must equal 8 in order to match the equation you were given. so we can see that we must have:\[2a=8\]therefore, dividing both sides by 2 we get:\[a=4\]

OpenStudy (asnaseer):

so now we know that:\[(x-4)^2\]will give us an expression with \(x^2-8x\) in it plus some constant. if we expand this we see this comes to:\[(x-4)^2=x^2-8x+16\]

OpenStudy (asnaseer):

now since your original expression did not contain a "+16", we need to add 16 to the right hand side of your original expression to balance the equations.

OpenStudy (asnaseer):

so we end up with:\[x^2-8x=-10\]therefore:\[(x-4)^2=-10+16=6\]

OpenStudy (asnaseer):

and we end up with something on the left-hand-side which is an exact square. this is called completing the square.

OpenStudy (asnaseer):

I hope you were able to follow the steps?

OpenStudy (anonymous):

Yes thanks so much!

OpenStudy (asnaseer):

yw